OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

see the steps.. use thies eq.\[\frac{ y-y1 }{ x-x1 }=\frac{ y1-y2 }{ x1-x2 }\] put here x1=-2,x2=5,y1=2,y2=1

OpenStudy (anonymous):

i said answers!!!!!!!!!!!

OpenStudy (anonymous):

7y-14=x+2

OpenStudy (anonymous):

First put it into Point-Slope Form....